OpenStudy (karinamtz0213):

√90 can be written as c√d , where d has no perfect square factors and c and d are both integers. Which of the following are equivalent to d-c: a. 1 b. 7 c. 13 d. 15 e. 19

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{90} = \sqrt{9*10}\] \[\large \sqrt{90} = \sqrt{9}*\sqrt{10}\] \[\large \sqrt{90} = 3\sqrt{10}\]

OpenStudy (karinamtz0213):

The answer you gave me is not a choice

jimthompson5910 (jim_thompson5910):

I know, compare it to c√d and tell me what c and d are (the values of each)

OpenStudy (karinamtz0213):

i don't know because how i posed the problem is exactly how it was given to me

jimthompson5910 (jim_thompson5910):

\[\large 3\sqrt{10} = c\sqrt{d}\]

jimthompson5910 (jim_thompson5910):

c = ?? d = ??

OpenStudy (karinamtz0213):

so then the answer must be 7.

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (karinamtz0213):

thanks

jimthompson5910 (jim_thompson5910):

np